Sally bought three chocolate bars and a pack of gum and paid $1.75. Jake bought two chocolate bars and four packs of gum and paid $2.00. Find the cost of a chocolate bar and the cost of a pack of gum.: Sally bought three chocolate bars and a pack of gum and paid $1.75. Jake bought two chocolate bars and four packs of gum and paid $2.00. Find the cost of a chocolate bar and the cost of a pack of gum.

Answer by ankor@dixie-net.com(12692)

Sally bought three chocolate bars and a pack of gum and paid $1.75. Jake bought two chocolate bars and four packs of gum and paid $2.00. Find the cost of a chocolate bar and the cost of a pack of gum.
:
Let x = cost of the chocolate bar
Let y = cost of the gum
:
Write two equations:
Sally: 3x + 1y = 1.75
Jake: 2x + 4y = 2.00
:
Multiply Sally's equation by 4, and subtract Jake's equation:
12x + 4y = 7.00
2x + 4y = 2.00
----------------- subtracting eliminates y
10x + 0y = 5.00
x = 5/10
x = .50 for chocolate bars
:
Find y using Sally's equation:
3(.50) + y = 1.75
1.50 + y = 1.75
y = 1.75 - 1.50
y = .25 for gum
:
:
Check solutions in Jake's equation:
2(.50) + 4(.25) =
1.00 + 1.00 = 2.00
:
:
How about that, did it make sense to you? Any questions?